OpenStudy (anonymous):

What is the slope of the line that passes through (-2, 2) and (-4, 5)? A.-3/2 B.3/2 C.-2/3 D.2/3

OpenStudy (anonymous):

-3/2

Directrix (directrix):

Slope of a line is the constant rate of change. m =( y1 - y2) all over( x1 - x2) where (x1,y1) and (x2, y2) are two points on the line. If (-2, 2) and (-4, 5) are points on the line and (x1,y1) = (-2,2) and (x2, y2) = (-4, 5). then what would m equal.

OpenStudy (anonymous):

\[\Large y-y_1=\frac{y_2-y_1}{x_2-x_1}(x-x_1)\] @Directrix I thought this was the formula O_O !!!

Directrix (directrix):

@Kreshnik --> That's the equation of the line. Question reads : What is the slope of the line

OpenStudy (anonymous):

ihhh... :F